### v1

@@ 1 Introduction
Artificial intelligence (AI), including Generative AI (GenAI) technology, has made significant advancements in various domains, including education, healthcare, and scientific research. GenAI models [10] can produce diverse content such as text, images, and videos using AI techniques. The integration of AI in education is increasing, albeit at a slower pace compared to other industries, yet its potential influence on education is substantial [11]. These technologies enable personalized educational materials, assessments, and tutoring [12]. Research [13] has indicated that incorporating AI into educational settings has the potential to enhance students’ autonomy in managing their own learning processes. Conversational AI, which conducts human-like conversations via text or audio based on LLMs, is widely accepted among students for task-oriented dialogues [12]. Tools like Google Assistant have improved EFL students’ communication skills and attitudes towards intelligent assistants in learning [14]. Chatbots providing feedback have effectively boosted vocabulary learning among Korean EFL primary students [15].
@@
The primary aim of this paper is to create an automated system to evaluate a candidate’s performance in the CEFR B2 English speaking test, removing the need for a human assessor and ensuring relevance in both global and Indian contexts. Secondary aims involve developing an automated system capable of identifying vocabulary and proficiency levels. Furthermore, this model should generate CEFR B2 vocabulary and sentences. We introduced EvalYaks, a set of six unique models, with four models dedicated to the primary aim and two to the secondary. EvalYaks is trained by instruction tuning Mistral Instruct 7B v0.2 using the Low Rank Adaptation (LoRA), a parameter efficient fine tuning (PEFT) method.
@@ 2 Cambridge’s B2 First Speaking Assessment Framework
We created EvalYaks utilizing the Cambridge framework for the B2 First Exam. This speaking assessment is a component of the comprehensive exam that evaluates all aspects of language skills, including listening, reading, writing, and speaking. The primary objective of this exam is to assess the English language proficiency of the candidate at the CEFR B2 level.
@@ 4.3 Performance metrics
where the acceptable accuracy for each part is indicated as $acceptable\_accuracy_{i}$, with $i=1,2,3,4$ referring to the different parts of the assessment.
@@ 6.1 Performance Analysis of standard LLMs without LoRA
Table 2: The average acceptable accuracy of leading LLMs without LoRA for CEFR B2 English speaking assessment using prompts without and with performance descriptors.
@@ 6.3 Performance analysis of EvalYaks Part 1-4 models
Figure 3: The average acceptable accuracy of leading LLMs without LoRA in comparison with EvalYaks part 1-4 models in CEFR B2 English speaking assessment with and without performance descriptors in the prompt.
Figure 4: The distribution of average acceptable accuracy of leading LLMs without LoRA in comparison with EvalYaks part 1-4 models in CEFR B2 English speaking assessment with and without performance descriptors in the prompt.
@@ 6.3.1 Analysis of EvalYaks Part 1-4 models without performance descriptors
Figure 5: The distribution of acceptable accuracy of leading LLMs without LoRA in comparison with EvalYaks part 1-4 models using prompts without performance descriptors.
@@ 6.3.2 Analysis of EvalYaks Part 1-4 models with performance descriptors
Figure 6: The distribution of acceptable accuracy of leading LLMs without LoRA in comparison with EvalYaks part 1-4 models using prompts with performance descriptors.
@@ 7 Conclusion
[…]
Acknowledgments
We appreciate the efforts of English language experts Binita Punwani, Cynthia Freeda B, and Varsha Kalyanakumar for validating and aligning the synthetic data with the scores based on Cambridge’s B2 First speaking exam performance descriptors.
@@ Acknowledgments
[…]
Author Contributions
All authors contributed to the study conception and design. Material preparation, data creation, training, and analysis were performed by Nicy Scaria and Silvester John Joseph Kennedy. The first draft of the manuscript was written by Nicy Scaria and Silvester John Joseph Kennedy together, and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
[…]
Avaliability of data and material
The instruction tuning datasets used for training EvalYaks can be accessed at https://github.com/Talking-Yak/EvalYaks. This repository contains only 30% of the data used for training. The entire data created for this research is available upon request. Please direct all inquiries to the corresponding author.
@@ Appendix C Evaluation prompts
The two sets of prompts that were used to evaluate the performance of leading LLMs without LoRA and ‘EvalYaks’ for part 1 are given below.

### v2

@@ 1 Introduction
Artificial intelligence (AI), including Generative AI (GenAI) technology, has made significant advancements in various domains, including education, healthcare, and scientific research. GenAI models [10] can produce diverse content such as text, images, and videos using AI techniques. The integration of AI in education is increasing, albeit at a slower pace compared to other industries, yet its potential influence on education is substantial [11]. These technologies enable personalized educational materials, assessments, and tutoring [12]. Research [13] has indicated that incorporating AI into educational settings has the potential to enhance students’ autonomy in managing their own learning processes. Conversational AI, which conducts human-like conversations via text or audio based on LLMs, is widely accepted among students for task-oriented dialogues [12]. Tools like Google Assistant have improved EFL students’ communication skills and attitudes towards intelligent assistants in learning [14]. The integration of LLMs with educational technology is being applied to a wide range of tasks, including automated grading of short answers [15, 16], essay scoring [17], generating questions automatically [18, 19, 20], creating multiple-choice questions [21, 22], and developing chatbots [23]. Chatbots providing feedback have effectively boosted vocabulary learning among Korean EFL primary students [24]. GPT-4 has shown better performance over human responses in open-ended assessments at various cognitive levels, highlighting its potential to support advanced educational evaluations [25].
@@
Figure 1: Cambridge’s CEFR B2 Speaking Assessment framework and it’s automated scoring using EvalYaks.
The primary aim of this paper is to create an automated system to evaluate and score a candidate’s performance in the CEFR B2 English speaking test, removing the need for a human assessor and ensuring relevance in both global and Indian contexts. The framework of CEFR B2 Speaking Assessment and the automated scoring system is given in Figure 1. Secondary aims involve developing an automated system capable of identifying vocabulary and proficiency levels. Furthermore, this model should generate CEFR B2 vocabulary and sentences. We introduced EvalYaks, a set of six unique models, with four models dedicated to the primary aim and two to the secondary. EvalYaks is trained by instruction tuning Mistral Instruct 7B v0.2 using the Low Rank Adaptation (LoRA), a parameter efficient fine tuning (PEFT) method.
@@ 2 Cambridge’s B2 First Speaking Assessment Framework
We created EvalYaks utilizing the Cambridge framework for the B2 First Exam. This speaking assessment is a component of the comprehensive exam that evaluates all aspects of language skills, including listening, reading, writing, and speaking. The primary objective of this exam is to assess the English language proficiency of the candidate at the CEFR B2 level. This framework is illustrated in Figure 1 and is explained in detail in the following subsections.
@@ 4.3 Performance metrics
where $acceptable\_accuracy_{i}$ is the acceptable accuracy for each part, with $i=1,2,3,4$ referring to the different parts of the assessment.
@@ 6.1 Performance Analysis of standard LLMs without LoRA
Table 2: The average acceptable accuracy of state-of-the-art LLMs without LoRA for CEFR B2 English speaking assessment using prompts without and with performance descriptors.
@@ 6.3 Performance analysis of EvalYaks Part 1-4 models
Figure 4: The average acceptable accuracy of state-of-the-art LLMs without LoRA in comparison with EvalYaks part 1-4 models in CEFR B2 English speaking assessment with and without performance descriptors in the prompt.
Figure 5: The distribution of average acceptable accuracy of state-of-the-art LLMs without LoRA in comparison with EvalYaks part 1-4 models in CEFR B2 English speaking assessment with and without performance descriptors in the prompt.
@@ 6.3.1 Analysis of EvalYaks Part 1-4 models without performance descriptors
Figure 6: The distribution of acceptable accuracy of state-of-the-art LLMs without LoRA in comparison with EvalYaks part 1-4 models using prompts without performance descriptors.
@@ 6.3.2 Analysis of EvalYaks Part 1-4 models with performance descriptors
Figure 7: The distribution of acceptable accuracy of state-of-the-art LLMs without LoRA in comparison with EvalYaks part 1-4 models using prompts with performance descriptors.
@@ 7 Conclusion
Acknowledgements
@@ Appendix C Evaluation prompts
The two sets of prompts that were used to evaluate the performance of state-of-the-art LLMs without LoRA and ‘EvalYaks’ for part 1 are given below.